OpenStudy (anonymous):

In a new suburb, it is estimated that the number of registered voters will be modeled by the function N(t) = 11 + 9 t^2 - t^3, where t is time in years and N is in thousands. i found n(6) = 119 The instantaneous rate of change of N at t = 6

OpenStudy (anonymous):

n(6) = 11 + 9*6^2 - 6^3 = 11 + 324 - 216 = 335 - 216 = 119

OpenStudy (anonymous):

i know i found that already

OpenStudy (anonymous):

wait is that the answer again?

OpenStudy (anonymous):

oh - im sorry - i thoght you wanted to confirm that

OpenStudy (anonymous):

u want instantaneous rate of change

OpenStudy (anonymous):

no i need to find The instantaneous rate of change of N at t = 6

OpenStudy (anonymous):

ok differentiate dN/dt = 18t - 3t^2 at t = 6 this is 18*6 - 3*36 = 108 - 108 = 0

OpenStudy (anonymous):

yes thank very much!

OpenStudy (anonymous):

thanx for the help

OpenStudy (anonymous):

yw